# Option-aware Temporally Abstracted Value for Offline Goal-Conditioned Reinforcement Learning

### v1

@@ Abstract
Offline goal-conditioned reinforcement learning (GCRL) offers a practical learning paradigm where goal-reaching policies are trained from abundant unlabeled (reward-free) datasets without additional environment interaction. However, offline GCRL still struggles with long-horizon tasks, even with recent advances that employ hierarchical policy structures, such as HIQL [34]. By identifying the root cause of this challenge, we observe the following insights: First, performance bottlenecks mainly stem from the high-level policy’s inability to generate appropriate subgoals. Second, when learning the high-level policy in the long-horizon regime, the sign of the advantage signal frequently becomes incorrect. Thus, we argue that improving the value function to produce a clear advantage signal for learning the high-level policy is essential. In this paper, we propose a simple yet effective solution: Option-aware Temporally Abstracted value learning, dubbed OTA, which incorporates temporal abstraction into the temporal-difference learning process. By modifying the value update to be option-aware, the proposed learning scheme contracts the effective horizon length, enabling better advantage estimates even in long-horizon regimes. We experimentally show that the high-level policy extracted using the OTA value function achieves strong performance on complex tasks from OGBench [33], a recently proposed offline GCRL benchmark, including maze navigation and visual robotic manipulation environments.
@@ 1 Introduction
Offline goal-conditioned reinforcement learning (GCRL) has emerged as a practical framework for real-world applications by leveraging pre-collected datasets to train goal-reaching policies without requiring additional environment interaction [24, 33]. However, learning an accurate goal-conditioned value function in long-horizon settings remains a major challenge: Naively training a goal-conditioned value function often leads to noisy estimates and erroneous policies [36, 21, 34]. To mitigate the learning of an erroneous policy, Hierarchical Implicit Q-Learning (HIQL) [34], one of the state-of-the-art methods, adopts a simple hierarchical structure in which a high-level policy predicts subgoals, and a low-level policy learns to execute actions toward those subgoals. In such a hierarchical structure, the separated policy extraction with same value function has different objectives for each policy. Though a hierarchical policy is still extracted from the noisy value function, both policies receive more reliable learning signals than when training a flat, non-hierarchical policy. However, despite reasonable performance gains of hierarchical methods in some long-horizon environments, a recent challenging benchmark [33] reveals that such a hierarchical policy still cannot solve more complex tasks, such as long-horizon robotic locomotion or robotic manipulation.
@@
To understand the failure in complex tasks more deeply, we raise the following question: Low-level policy vs. high-level policy: Which is the Bottleneck of HIQL? To answer this question, we analyze the hierarchical policy in failure cases by generating oracle subgoals for the low-level policy. Interestingly, we observe that the low-level policy achieves subgoals with high accuracy, indicating that the failure stems from the inability of the high-level policy to generate appropriate subgoals. The main reason for the limited performance of the high-level policy is that the noisy value function still cannot provide useful learning signals for training the high-level policy in long-horizon scenarios.
@@
Based on the phenomenon that the high-level policy eventually failed to obtain useful learning signals from the value function, the main cause of noisy learning signals is the order inconsistency of the learned value function in the long-horizon setting. Our analysis reveals that when the distance between the state and the goal exceeds a certain temporal horizon, a critical issue arises with the advantage signal. In the long-horizon regime, it is frequently observed that the sign of the advantage signal is incorrect, causing erroneous regression weights for learning the high-level policy. Considering the issue with the value function, we argue that designing a value function that can produce a clear advantage signal for extracting the high-level policy is necessary.
@@
Motivated by the observation that the low-level policy performs remarkably well at reaching short-horizon subgoals, we propose a simple yet effective value function learning scheme for high-level policy extraction by reducing the horizon between the state and the goal. Leveraging the notion of option [46], also known as macro actions consisting of a sequence of primitive actions, we redesign the value learning scheme to be option-aware. Specifically, by updating the value over a sequence of primitive actions, the effective horizon between the state and the goal is significantly reduced compared to the primitive action-aware value learning scheme [22]. In our experiment, we show that our value learning scheme effectively mitigates the errors of the value function in the long-horizon regime. Furthermore, we evaluate our approach in various tasks, including maze and robotic visual manipulation environments from OGBench [33], and empirically demonstrate that using our value function to extract the high-level policy yields superior performance on long-horizon tasks compared to baselines.
@@
  • Through analysis of the failure cases of hierarchical policies, we identify that the failures stem from the inability of the high-level policy to generate appropriate subgoals. Furthermore, we observe that the value function used for extracting the high-level policy has significant errors when the distance between the state and the goal is large.
  • To tackle this problem, we propose Option-aware Temporally Abstracted (OTA) value learning, which produces a reduced horizon compared to the conventional value learning objective [22].
  • In our experiments, we demonstrate that, despite long state-to-goal horizons, our value function yields significantly lower errors, and the hierarchical policy extracted using our value function successfully solves complex maze and robotic manipulation tasks.
@@ 2 Related Work
GCRL. GCRL aims to train goal-conditioned policies to reach arbitrary goal states from given initial states, rather than optimizing for a single, fixed task [43, 26]. Our work focuses specifically on offline GCRL [4, 27, 53, 34, 44, 33], in which goal-conditioned policies are learned entirely from pre-collected datasets without further environment interaction. Due to the sparse rewards in goal-reaching tasks, offline GCRL has relied on hindsight data relabeling [1, 41, 56], and more recently, imitation learning and value-based methods have been explored to better leverage suboptimal datasets [6, 12, 53, 13]. In these works, the value function is typically learned through temporal-difference (TD) methods [22, 35], or through alternative techniques such as state-occupancy matching [27, 7], contrastive learning [28, 9, 25] and quasimetric learning [51]. However, whether the value functions can effectively generalize to long-horizon tasks remains an open question [33].
@@
Hierarchical RL. Achieving long-horizon goals remains a fundamental challenge in GCRL [38, 36, 21, 34]. To address this, hierarchical RL methods have adopted either graph-based planning in the state space [8, 17, 55, 21, 54, 20] or waypoint-based subgoal generation [5, 23, 48, 29, 18, 14, 32, 31, 19, 3, 34, 52]. However, graph-based planning methods incur high computational overhead and architectural complexity. Waypoint-based approaches also face challenges in generating effective subgoals in long-horizon settings, due to inaccurate value estimates when the state is far from the goal.
@@
Option framework. To enhance the planning capabilities of an agent over long time horizons, one effective approach is to leverage temporal abstraction through the option framework, which involves learning sub-policies known as options [15, 46, 42, 36]. In this framework, options serve as temporally extended actions that enable planning across multiple time scales. After establishing the theoretical connection between the option framework and semi-Markov decision processes [46], research has progressed toward end-to-end option learning [40, 45, 2, 47] and automatic option discovery [39, 16]. Our method is closely related to HIQL [34], which trains a high-level policy to generate subgoals. However, unlike HIQL, our approach leverages options defined in offline datasets to effectively reduce the planning horizon during value function training. As a result, our high-level policy can generate subgoals over longer temporal horizons without relying on explicit option learning or option discovery.
@@ 3 Preliminaries
Hierarchical Implicit Q-Learning (HIQL).   In GCRL, accurately estimating the value function for distant goals is the main challenge in solving complex long-horizon tasks [17, 21, 34]. To address this issue, HIQL [34] proposed a hierarchical policy structure that utilizes a value function learned with IQL [22]. This hierarchical design enables the agent to produce effective actions even when value estimates for distant goals are noisy or unreliable. More specifically, HIQL trains a goal-conditioned state-value function $V$ with the following loss:
@@
where the expectile loss is defined as $L_{2}^{\tau}(u)=|\tau-\mathbf{1}(u<0)|u^{2}$, with $\tau>0.5$, and $\bar{V}$ denotes the target $V$ network. Following prior works [1, 8, 3, 51, 34, 52], we adopt the sparse reward $r(s,g)=-\mathbf{1}\{s\neq g\}$. Under this reward, the optimal value $|V^{\star}(s,g)|$ corresponds to the discounted temporal distance, i.e., a discounted measure of the minimum number of environment steps required to reach the goal $g$ from state $s$. HIQL separates policy extraction into two levels: a high-level policy $\pi^{h}(s_{t+k}|s_{t},g)$ generates a $k$-step subgoal to guide progress toward the goal, while a low-level policy $\pi^{\ell}(a_{t}|s_{t},s_{t+k})$ produces primitive actions to reach the subgoal. Both policies are learned using advantage-weighted regression (AWR) [49, 37, 30] with the following objective:
@@
where $\beta^{h}$ and $\beta^{l}$ are inverse temperature parameters, $A^{h}(s_{t},s_{t+k},g)=V^{h}(s_{t+k},g)-V^{h}(s_{t},g)$ denotes the high-level policy advantage, and $A^{\ell}(s_{t},s_{t+1},s_{t+k})=V^{\ell}(s_{t+1},s_{t+k})-V^{\ell}(s_{t},s_{t+k})$ denotes the low-level policy advantage. HIQL uses a single goal-conditioned value function $V$, which is shared between both $\pi^{h}$ and $\pi^{\ell}$ (i.e., $V^{h}=V^{\ell}=V$). However, despite this design, HIQL still struggles with long-horizon, complex tasks, as shown in the recent offline GCRL benchmark, OGBench [33].
@@ 4 Motivation
In this subsection, we investigate the failure cases of HIQL in long-horizon scenarios by identifying whether the main performance bottleneck is in the low-level policy or the high-level policy. To examine this, we fix the low-level policy $\pi^{\ell}$ and replace the high-level policy $\pi^{h}$ with an oracle policy, $\pi^{h}_{\text{oracle}}$, which always provides optimal subgoals reachable within a short horizon. [^1] Specifically, $\pi^{h}_{\text{oracle}}$ provides as a subgoal the center of an adjacent maze cell that lies on the shortest path from the current state to the goal. We refer to this variant as $\text{HIQL}^{\text{OS}}$, and pose the following hypothesis: if $\text{HIQL}^{\text{OS}}$ still fails in long-horizon tasks, then the low-level policy struggles to reach short-horizon subgoals. Conversely, if it achieves a high success rate, the main problem lies in the high-level policy.
Figure 1 shows the results of HIQL and $\text{HIQL}^{\text{OS}}$ on eight challenging maze navigation tasks from OGBench [33]. HIQL achieves less than 20% success rate on many tasks, indicating that HIQL significantly fails to solve the long-horizon tasks. In contrast, we note that $\text{HIQL}^{\text{OS}}$ achieves a much higher success rate around 90%. These results indicate that, while the low-level policy generalizes well in short-horizon settings when provided with accurate subgoals, the inaccuracy of the high-level policy is the primary cause of HIQL’s failure in long-horizon scenarios.
@@
We identify two potential issues in Equation (2) that may underlie the failure of high-level policy learning: 1) an inadequate policy extraction scheme (i.e., the regression component in Equation (2)), and 2) an inaccurately learned value function (i.e., the advantage term in Equation (2)). Since the same policy extraction scheme enables successful low-level policy learning, we do not consider it to be the primary cause of failure. This suggests that the inaccurate value function used in the high-level policy advantage term may be the key contributor to the failure. In particular, as the distance between $s_{t}$ and $g$ increases, the value estimates become increasingly erroneous, leading to an imprecise evaluation of the high-level advantage. Although HIQL attempts to mitigate the noise in estimating the long-horizon value $V^{h}$ through its hierarchical structure, it is possible that the high-level advantage may still be significantly erroneous. In the following subsection, we carefully analyze how such errors in estimating $V^{h}$ adversely affect high-level policy learning.
@@ 4.2 Order Inconsistency of the Learned Value Function in the Long-Horizon Setting
(Order consistency) Let $\tau^{\star}=(s_{0},s_{1},\ldots,s_{T}=g)$ denote the optimal trajectory induced by the optimal policy $\pi^{\star}(\cdot\mid s,g)$, from the initial state $s_{0}$ to the goal $g$, and let $V$ be a learned value function. Given $s_{i},s_{j}\in\tau^{*}$ with $j>i$, we say that $V$ satisfies order consistency with respect to $(s_{i},s_{j},g)$ if and only if $V(s_{j},g)>V(s_{i},g)$.
@@
Consider an optimal trajectory $\tau^{\star}=\{s_{0},s_{1},\dots,s_{T}\}$, generated by an oracle policy. Along this trajectory, the optimal value function is increasing, such that $V^{\star}(s_{j},g)>V^{\star}(s_{i},g)$ for all $j>i$. Thus, value order consistency refers to the alignment between the order induced by $V^{h}$ and that induced by $V^{\star}$. We argue that achieving the order consistency between $V(s_{t},g)$ and $V(s_{t+k},g)$ is critical, as sign mismatches can invert the high-level advantage signal $A^{h}$ and hinder the learning of an appropriate high-level policy.
@@
To check whether the learned $V^{h}$ of HIQL achieves the order consistency or not, we have collected optimal trajectories across four different long-horizon tasks with specified goals using near-optimal policies, as illustrated in Figure 2. The trajectory lengths varied from 250 to 2000 steps. For each state $s_{t}$ in the trajectory, we then visualize the learned value $V^{h}(s_{t},g)$ alongside the optimal value function, computed as $V^{\star}(s_{t},g)=-\left(1-\gamma^{d^{\star}(s_{t},g)}\right)/\left(1-\gamma\right)$, in which $d^{\star}(s_{t},g)$ denotes the temporal distance between $s_{t}$ and $g$. Since the value decays exponentially as the distance to the goal increases due to the discount factor $\gamma$, it becomes difficult to visually interpret the relative differences in value. Hence, we convert $V^{h}(s,g)$ into estimated temporal distances using the following equation: $d^{h}(s,g)=\log\left(1+(1-\gamma)V^{h}(s,g)\right)/\log\gamma$. In this form, the condition for value order consistency is equivalent to $d^{h}(s_{i},g)>d^{h}(s_{j},g)$, where $j>i$.
@@
As shown in Figure 2, we note that $V^{h}$ closely matches $V^{\star}$ when the state is near the goal (i.e., $d^{\star}(s,g)\approx 0$). This alignment explains the strong performance of the low-level policy presented in Figure 1. However, when the state-goal distance exceeds a certain temporal horizon, the value order inconsistency frequently arises between $V^{h}(s_{t},g)$ and $V^{h}(s_{t+k},g)$ due to the non-monotonicity of the learned $V^{h}$. [^2] The hyperparameter $k$ in HIQL is chosen based on the characteristics of the environments and datasets. In Figure 2, $k=25$ for the AntMaze task and $k=100$ for the HumanoidMaze task. This is due to the well-known fact that the learning target for the value in Equation (1) becomes noisier as the horizon becomes longer, and shows why the use of $V^{h}$ becomes less effective in high-level policy learning for long-horizon setting.
@@
Motivated by the observation that $V^{h}$ aligns well with $V^{\star}$ and achieves order consistency in short-horizon settings, in the next section, we propose a simple yet effective solution based on temporal abstraction [46]. This approach enables high-level value function learning to provide appropriate advantage signals, even when $d^{\star}(s,g)$ is large.

### v2

@@ Abstract
Offline goal-conditioned reinforcement learning (GCRL) offers a practical learning paradigm in which goal-reaching policies are trained from abundant state–action trajectory datasets without additional environment interaction. However, offline GCRL still struggles with long-horizon tasks, even with recent advances that employ hierarchical policy structures, such as HIQL (HIQL)Park2023 . Identifying the root cause of this challenge, we observe the following insight. Firstly, performance bottlenecks mainly stem from the high-level policy’s inability to generate appropriate subgoals. Secondly, when learning the high-level policy in the long-horizon regime, the sign of the advantage estimate frequently becomes incorrect. Thus, we argue that improving the value function to produce a clear advantage estimate for learning the high-level policy is essential. In this paper, we propose a simple yet effective solution: Option-aware Temporally Abstracted value learning, dubbed OTA, which incorporates temporal abstraction into the temporal-difference learning process. By modifying the value update to be option-aware, our approach contracts the effective horizon length, enabling better advantage estimates even in long-horizon regimes. We experimentally show that the high-level policy learned using the OTA value function achieves strong performance on complex tasks from OGBench (OGBench)Park2025 , a recently proposed offline GCRL benchmark, including maze navigation and visual robotic manipulation environments. Our code is available at https://github.com/ota-v/ota-v
@@ 1 Introduction
Offline goal-conditioned reinforcement learning (GCRL) has emerged as a practical framework for real-world applications by leveraging pre-collected datasets to train goal-reaching policies without requiring additional environment interaction (offlineRL)levine2020 ; (OGBench)Park2025 . However, learning an accurate goal-conditioned value function in long-horizon settings remains a major challenge, as naively training the value function often leads to noisy estimates and erroneous policies (HRLsurvey)Pateria2021 ; (HIGL)Kim2021 ; (HIQL)Park2023 . To mitigate the learning of an erroneous policy, Hierarchical Implicit Q-Learning (HIQL) (HIQL)Park2023 , one of the state-of-the-art methods, adopts a simple hierarchical structure in which a high-level policy predicts subgoals, and a low-level policy learns to execute actions toward those subgoals. Though a hierarchical policy is still learned from the noisy value function, both policies receive more reliable learning signals than when training a flat, non-hierarchical policy. However, despite reasonable performance gains of hierarchical methods in some long-horizon environments, a recent challenging benchmark (OGBench)Park2025 reveals that such a hierarchical policy still cannot solve more complex tasks, such as long-horizon robotic locomotion or robotic manipulation.
@@
To understand the failure in complex tasks more deeply, we raise the following question: Low-level policy vs. high-level policy: which is the bottleneck of HIQL? To answer this question, we analyze the hierarchical policy in failure cases by generating oracle subgoals for the low-level policy. Interestingly, we observe that the low-level policy achieves these subgoals with high accuracy, indicating that the failure stems from the inability of the high-level policy to generate appropriate subgoals. The limited performance primarily results from a noisy value function, which fails to provide sufficiently informative learning signals for effectively training the high-level policy in long-horizon scenarios.
@@
Based on the phenomenon that the high-level policy eventually failed to extract meaningful learning signals from the value function, we identify the primary cause of these noisy signals as the order inconsistency of the learned value function in the long-horizon setting. Our analysis reveals that when the distance between the state and the goal exceeds a certain temporal horizon, the sign of the advantage estimate is incorrect, causing erroneous regression weights for learning the high-level policy. Considering the issue with the value function, we argue that designing a value function that can produce a clear advantage estimate for learning the high-level policy is necessary.
@@
Motivated by the observation that the low-level policy performs remarkably well at reaching short-horizon subgoals, we propose a simple yet effective value function learning scheme for high-level policy learning that reduces the horizon between the state and the goal. Specifically, we leverage the notion of option (option)Sutton1999 , a temporally-extended course of action, by updating the value over sequences of primitive actions. This option-aware value learning substantially shortens the effective horizon compared to primitive action-aware value learning (IQL)Kostrikov , mitigating errors in long-horizon value estimation. We evaluate our approach on maze and robotic visual manipulation tasks from OGBench (OGBench)Park2025 , and empirically show that using our value function enables the high-level policy to achieve superior performance on long-horizon tasks.
@@
  • Through analysis of the failure cases of hierarchical policies, we identify that the failures stem from the inability of the high-level policy to generate appropriate subgoals. Furthermore, we observe that the value function used for high-level policy learning has significant errors when the distance between the state and the goal is large.
  • To tackle this problem, we propose Option-aware Temporally Abstracted (OTA) value learning, which reduces the effective horizon compared to the conventional value learning objective (IQL)Kostrikov .
  • Our experiments show that, even across long state-to-goal horizons, our value function achieves significantly lower errors, enabling the hierarchical policy to successfully solve complex maze and robotic manipulation tasks.
@@ 2 Related Work
GCRL. GCRL aims to train goal-conditioned policies to reach arbitrary goal states from given initial states, rather than optimizing for a single, fixed task (UVFA)Schaul2015 ; (GCRLsurey)Liu2022 . Our work focuses specifically on offline GCRL (AM)Chebotar2021 ; (GoFAR)Ma2022 ; (WGCSL)Yang2022 ; (HIQL)Park2023 ; (SM)Sikchi2024 ; (OGBench)Park2025 , in which goal-conditioned policies are learned entirely from pre-collected datasets without further environment interaction. Due to the sparse rewards in goal-reaching tasks, offline GCRL has relied on hindsight data relabeling (HER)Andrychowicz2017 ; (HGG)Ren2019 ; (GOALIVE)Zheng2024 , and more recently, imitation learning and value-based methods have been explored to better leverage suboptimal datasets (goalGAIL)Ding2019 ; (GCSL)Ghosh2019 ; (WGCSL)Yang2022 ; (GCPO)gong2024 . In these works, the value function is typically learned through temporal-difference (TD) methods (IQL)Kostrikov ; (HILP)Park2024 , or through alternative techniques such as state-occupancy matching (GoFAR)Ma2022 ; (adversarial)durugkar2021 , contrastive learning (VIP)Ma2022 ; (CRL)Eysenbach2022 ; (single)Liu2025 , and quasimetric learning (QRL)Wang2023 . However, whether the value functions can effectively generalize to long-horizon tasks remains an open question (OGBench)Park2025 .
@@
Hierarchical RL. Achieving long-horizon goals remains a fundamental challenge in GCRL (TA)Precup2000 ; (HRLsurvey)Pateria2021 ; (HIGL)Kim2021 ; (HIQL)Park2023 . To address this, hierarchical RL methods have adopted either graph-based planning in the state space (SoRB)eysenbach2019 ; (MSS)Huang2019 ; (HRAC)Zhang2020 ; (HIGL)Kim2021 ; (L3P)Zhang2021 ; (PIG)kim2023 or waypoint-based subgoal generation (Feudal)Dayan1992 ; (h-DQN)Kulkarni2016 ; (FuNs)Vezhnevets2017 ; (HIRO)nachum2018 ; (L-Abstraction)Jiang2019 ; (HiTS)gurtler2021 ; (LEAP)nasiriany2019 ; (HVF)Nair2020 ; (SGT)Jurgenson2020 ; (RIS)chane2021 ; (HIQL)Park2023 ; (MSCP)Wu2024 . However, graph-based planning methods incur high computational overhead and architectural complexity. Waypoint-based approaches also face challenges in generating effective subgoals in long-horizon settings, due to inaccurate value estimates when the state is far from the goal.
@@
Option framework. To enhance the planning capabilities of an agent over long time horizons, one effective approach is to leverage temporal abstraction through the option framework, which involves learning sub-policies known as options (HS)Hauskrecht1998 ; (option)Sutton1999 ; (role)Riemer2020 ; (HRLsurvey)Pateria2021 . In this framework, options serve as temporally extended actions that enable planning across multiple time scales. After establishing the theoretical connection between the option framework and semi-Markov decision processes (option)Sutton1999 , research has progressed toward end-to-end option learning (macroactions)Randlov1998 ; (OptionLearning)Stolle2002 ; (OptionCritic)Bacon2017 ; (STRAW)Vezhnevets2016 and automatic option discovery (successoroptoin)Ramesh2019 ; (OptionZero)Huang2025 . Our method is closely related to HIQL (HIQL)Park2023 , which trains a high-level policy to generate subgoals. However, unlike HIQL, our approach leverages options defined in offline datasets to effectively reduce the planning horizon during value function training. As a result, our high-level policy can generate subgoals over longer temporal horizons without relying on explicit option learning or option discovery.
@@ 3 Preliminaries
Hierarchical Implicit Q-Learning (HIQL). In GCRL, accurately estimating the value function for distant goals is the main challenge in solving complex long-horizon tasks (MSS)Huang2019 ; (HIGL)Kim2021 ; (HIQL)Park2023 . To address this issue, HIQL (HIQL)Park2023 proposed a hierarchical policy structure that utilizes a value function learned with IQL (IQL)Kostrikov . This hierarchical design enables the agent to produce effective actions even when value estimates for distant goals are noisy or unreliable. More specifically, HIQL trains a goal-conditioned state-value function $V$ with the following loss:
@@
where the expectile loss is defined as $L_{2}^{\tau}(u)=|\tau-\mathbf{1}(u<0)|u^{2}$, with $\tau>0.5$, and $\bar{V}$ denotes the target $V$ network. [^1] Note that since the inherent over-estimation problem of IQL, in this paper, we assume that the environment dynamics is deterministic. Following prior works (HER)Andrychowicz2017 ; (SoRB)eysenbach2019 ; (RIS)chane2021 ; (QRL)Wang2023 ; (HIQL)Park2023 ; (MSCP)Wu2024 , we adopt the sparse reward $r(s,g)=-\mathbf{1}\{s\neq g\}$. Under this reward, the optimal value $|V^{\star}(s,g)|$ corresponds to the discounted temporal distance, i.e., a discounted measure of the minimum number of environment steps required to reach the goal $g$ from state $s$. HIQL separates policy extraction [^2] Policy extraction refers to learning a policy from a learned value function, emphasizing the separation between value learning and policy learning. into two levels: a high-level policy $\pi^{h}(s_{t+k}|s_{t},g)$ generates a $k$-step subgoal to guide progress toward the goal, while a low-level policy $\pi^{\ell}(a_{t}|s_{t},s_{t+k})$ produces primitive actions to reach the subgoal. Both policies are extracted using advantage-weighted regression (AWR) (MARWIL)Wang2018 ; (AWR)Peng2019 ; (AWAC)Nair2020 with the following objective:
@@
where $\beta^{h}$ and $\beta^{l}$ are inverse temperature parameters, $A^{h}(s_{t},s_{t+k},g)=V^{h}(s_{t+k},g)-V^{h}(s_{t},g)$ denotes the high-level policy advantage, and $A^{\ell}(s_{t},s_{t+1},s_{t+k})=V^{\ell}(s_{t+1},s_{t+k})-V^{\ell}(s_{t},s_{t+k})$ denotes the low-level policy advantage. HIQL uses a single goal-conditioned value function $V$, which is shared between both $\pi^{h}$ and $\pi^{\ell}$ (i.e., $V^{h}=V^{\ell}=V$). However, despite this design, HIQL still struggles with long-horizon, complex tasks, as shown in the recent offline GCRL benchmark, OGBench (OGBench)Park2025 .
@@ 4 Motivation
We investigate the failure cases of HIQL in long-horizon scenarios by identifying whether the main performance bottleneck is in the low-level policy or the high-level policy. To examine this, we fix the low-level policy $\pi^{\ell}$ and replace the high-level policy $\pi^{h}$ with an oracle policy $\pi^{h}_{\text{oracle}}$, which always provides optimal subgoals reachable within a short horizon. [^3] Specifically, $\pi^{h}_{\text{oracle}}$ provides as a subgoal the center of an adjacent maze cell that lies on the shortest path from the current state to the goal. We refer to this variant as $\text{HIQL}^{\text{OS}}$, and pose the following hypothesis: if $\text{HIQL}^{\text{OS}}$ still fails in long-horizon tasks, then the low-level policy struggles to reach short-horizon subgoals. Conversely, if it achieves a high success rate, the main problem lies in the high-level policy.
Figure 1 shows the results of HIQL and $\text{HIQL}^{\text{OS}}$ on eight challenging maze navigation tasks from OGBench (OGBench)Park2025 . HIQL achieves less than 20% success rate on many tasks, indicating that HIQL significantly fails to solve the long-horizon tasks. In contrast, we note that $\text{HIQL}^{\text{OS}}$ achieves a much higher success rate around 90%. These results indicate that, although the low-level policy generalizes well in short-horizon settings when provided with accurate subgoals, the primary failure of HIQL in long-horizon scenarios stems from inaccuracies in the high-level policy.
@@
We identify two potential issues in Equation (2) that may underlie the failure of high-level policy learning: (1) an inadequate policy extraction scheme (i.e., the regression component in Equation (2)), and (2) an inaccurately learned value function (i.e., the advantage term in Equation (2)). Since the same policy extraction scheme enables successful low-level policy learning, we do not consider it to be the primary cause of failure. This suggests that the inaccurate value function used in the high-level policy advantage term may be the key contributor to the failure. In particular, as the distance between $s_{t}$ and $g$ increases, the value estimates become increasingly erroneous, leading to an imprecise evaluation of the high-level advantage. Although HIQL attempts to mitigate the noise in estimating the long-horizon value $V^{h}$ through its hierarchical structure, the high-level advantage may still be substantially erroneous. In the following subsection, we carefully analyze how such errors in estimating $V^{h}$ adversely affect high-level policy learning.
@@ 4.2 Order Inconsistency of the Learned Value Function in the Long-Horizon Setting
(Order consistency) Assume that the environment is deterministic. Let $\tau^{\star}=(s_{0},s_{1},\ldots,s_{T}=g)$ denote the optimal trajectory induced by the optimal policy $\pi^{\star}(\cdot\mid s,g)$, from the initial state $s_{0}$ to the goal $g$, and let $V$ be a learned value function. Given $s_{i},s_{j}\in\tau^{*}$ with $j>i$, we say that $V$ satisfies order consistency with respect to $(s_{i},s_{j},g)$ if and only if $V(s_{j},g)>V(s_{i},g)$.
@@
Consider an optimal trajectory $\tau^{\star}=\{s_{0},s_{1},\dots,s_{T}\}$, generated by an oracle policy. Along this trajectory, the optimal value function is increasing, such that $V^{\star}(s_{j},g)>V^{\star}(s_{i},g)$ for all $j>i$. Thus, value order consistency refers to the alignment between the order induced by $V^{h}$ and that induced by $V^{\star}$. We argue that achieving the order consistency between $V(s_{t},g)$ and $V(s_{t+k},g)$ is critical, as sign mismatches can invert the high-level advantage estimate $A^{h}$ and hinder the learning of an appropriate high-level policy. With large $k$ values (e.g., $25$ in AntMaze $100$ in HumanoidMaze), the sign mismatch of the advantage estimate can lead to significant performance degradation. When the advantage sign is incorrect, the magnitude of regression weights (which is the exponentiated advantage) is drastically increased or decreased, leading to improper subgoal regression for high-level policy. For example, if the range of an advantage is [1,1][-1,1] with $\beta^{h}=3$, the regression weights vary from $e^{-3}\approx 0.05$ to $e^{3}\approx 20$, indicating that a sign flip can significantly change the weight magnitude.
@@
To check whether the learned $V^{h}$ of HIQL achieves the order consistency or not, we collected optimal trajectories for four different long-horizon tasks with specified goals using near-optimal policies, as illustrated in Figure 2. The trajectory lengths varied from 250 to 2000 steps. For each state $s_{t}$ in the trajectory, we then visualize the learned value $V^{h}(s_{t},g)$ alongside the optimal value function, computed as $V^{\star}(s_{t},g)=-\left(1-\gamma^{d^{\star}(s_{t},g)}\right)/\left(1-\gamma\right)$, in which $d^{\star}(s_{t},g)$ denotes the temporal distance between $s_{t}$ and $g$. Since the value decays exponentially as the distance to the goal increases due to the discount factor $\gamma$, directly comparing relative values is visually challenging. Hence, we transform $V^{h}(s,g)$ into estimated temporal distances using the following equation: $d^{h}(s,g)=\log\left(1+(1-\gamma)V^{h}(s,g)\right)/\log\gamma$. In this form, the criterion for value order consistency becomes $d^{h}(s_{i},g)>d^{h}(s_{j},g)$, where $j>i$.
@@
As shown in Figure 2, we note that $V^{h}$ closely matches $V^{\star}$ when the state is near the goal (i.e., $d^{\star}(s,g)\approx 0$). This alignment explains the strong performance of the low-level policy presented in Figure 1. However, when the state-goal distance exceeds a certain temporal horizon, the value order inconsistency frequently arises between $V^{h}(s_{t},g)$ and $V^{h}(s_{t+k},g)$ due to the non-monotonicity of the learned $V^{h}$. [^4] The hyperparameter $k$ in HIQL is chosen based on the characteristics of the environments and datasets. In Figure 2, $k=25$ for the AntMaze task and $k=100$ for the HumanoidMaze task. This is due to the well-known fact that the learning target for the value in Equation (1) becomes noisier as the horizon becomes longer, and shows why the use of $V^{h}$ becomes less effective in high-level policy learning for long-horizon settings.
@@
Motivated by the observation that $V^{h}$ aligns well with $V^{\star}$ and achieves order consistency in short-horizon settings, we propose a simple yet effective solution based on temporal abstraction (option)Sutton1999 . This approach enables high-level value function learning to provide appropriate advantage estimates, even when $d^{\star}(s,g)$ is large.